OpenStudy (anonymous):

HELP! The following function shows the relationship between the selling prices, and profit P(s), in dollars, for a company: P(s) = -20s2 + 1,400s - 12,000 Which statement best describes the intervals where the company's profit increases, decreases, or records a maximum? a) It is least when the selling price is $30. b) It is greatest when the selling price is $30. c) It decreases when the selling price increases from $10 to $35. d) It increases when the selling price increases from $10 to $35. Explanation for answer would be greatly appreciated. Thanks.

OpenStudy (anaise):

hi

OpenStudy (solomonzelman):

Calc 1?

OpenStudy (solomonzelman):

Or just a vertex problem?

OpenStudy (anonymous):

I wish...but nah Algebra I

OpenStudy (solomonzelman):

Well, we don't need calculus really, since it is a parabola..

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle p(s)=-20s^2+1400s-12000 }\)

OpenStudy (solomonzelman):

Can you factor the left side out of -20?

OpenStudy (anonymous):

Sure, one sec.

OpenStudy (anonymous):

-20(s^2-700s+600)

OpenStudy (solomonzelman):

wait the middle

OpenStudy (solomonzelman):

middle term is wrong.

OpenStudy (anonymous):

Sorry, -20 (s^2 + 70s + 600)

OpenStudy (anonymous):

Wait isn't that wrong?

OpenStudy (anonymous):

Oh man i factored 14000 instead, one sec

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle -20 (s^2 - 70s + 600) }\)

OpenStudy (anonymous):

-20 (s-10) (s-60)

OpenStudy (solomonzelman):

no need for complete factorization.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle p(s)=-20 (s^2 - 70s + 600) }\) is what you need at this point.

OpenStudy (anonymous):

Okay i got that

OpenStudy (solomonzelman):

you need to complete the square... What would you add \(\color{#000000 }{ \displaystyle s^2 - 70s + ~? }\) to make a perfect square trinomial?

OpenStudy (solomonzelman):

Note that, \(\color{#000000 }{ \displaystyle x^2+2ax+a^2\color{grey}{~~~~~=(x+a)^2}}\) is something that you want to obtain

OpenStudy (solomonzelman):

so the 2a corresponds (in our case), to -70, right?

OpenStudy (anonymous):

It's not a perfect square, right?

OpenStudy (solomonzelman):

the x²+2ax+a² is a perf. sq.

OpenStudy (solomonzelman):

your polynomial inside the parenthesis is not,,,

OpenStudy (anonymous):

Yea, in the parenthesis, the first term and last term are perfect squares, but the middle one isn't, right?

OpenStudy (solomonzelman):

but if you tell me the missing number for \(\color{#000000 }{ \displaystyle s^2 - 70s +~{\rm what?} }\) then I can show you a trick.

OpenStudy (solomonzelman):

"what?" will make the s²-70s a perfect square trinomial when added?

OpenStudy (anonymous):

I don't know

OpenStudy (solomonzelman):

\(s^2+2as+a^2\) is the form you want

OpenStudy (solomonzelman):

you have \(s^2-70s\)

OpenStudy (solomonzelman):

So -70s, corresponds to the "2as" peace, right?

OpenStudy (solomonzelman):

So, what is the "a" in our case?

OpenStudy (anonymous):

-35?

OpenStudy (solomonzelman):

yup

OpenStudy (solomonzelman):

and "a²" is what?

OpenStudy (anonymous):

YAY!

OpenStudy (solomonzelman):

ok, next question i asked...

OpenStudy (anonymous):

(-35)^2

OpenStudy (solomonzelman):

yes, and that would be?

OpenStudy (anonymous):

1225

OpenStudy (solomonzelman):

yup

OpenStudy (solomonzelman):

So, I will show you the trick now...

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle p(s)=-20 (s^2 - 70s + 600) }\) You already have 600 in parenthesis, and you need to make that a 1225. (For this to be a perfect square trinomial) \(\color{#000000 }{ \displaystyle 600+x=1225 }\) \(\color{#000000 }{ \displaystyle x=625 }\) But, you can't just add 625, you will have to use something that I refer to as the "magic zero", and you will see what I am talking about NOW... \(\color{#000000 }{ \displaystyle p(s)=-20 (s^2 - 70s + 600+625-625) }\) Now, we would like to get rid of -625 in parenthesis, but we can't just erase it. We will take it out, by multiplying times -20.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle p(s)=-20 (s^2 - 70s + 600+625) +(-20)(-625)}\) \(\color{#000000 }{ \displaystyle p(s)=-20 (s^2 - 70s + 1225) +1250}\) and recall the form, \(s^2+2as+a^2=(s+a)^2\) We have already clarified that: a=-35 a²=1225 2a=-70 \(\color{#000000 }{ \displaystyle p(s)=-20 (s^2 - 2(35)s +(-35)^2) +1250}\)

OpenStudy (solomonzelman):

And we apply the form now! \(\color{#000000 }{ \displaystyle p(s)=-20 (s - 35)^2 +1250}\)

OpenStudy (solomonzelman):

when you digest this info, let me know...

OpenStudy (anonymous):

One moment, still digesting

OpenStudy (anonymous):

I got it. What's next?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle p(s)=-20 (s - 35)^2 +1250}\) can you tell me if this parabola opens up or down (and why)?

OpenStudy (anonymous):

down? i think..

OpenStudy (solomonzelman):

yes, and why?

OpenStudy (anonymous):

negative coefficient?

OpenStudy (solomonzelman):

yes, the leading coefficient is negative... fabulous!

OpenStudy (solomonzelman):

Since the parabola opens down, the vertex is the maximum point of the parabola. (The other points are all lower)

OpenStudy (solomonzelman):

So you need to find the vertex, and for this there is a rule...

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y(x)=a(x-k)^2+h }\) will have a vertex of \(\color{#000000 }{ \displaystyle (h,k) }\)

OpenStudy (anonymous):

(1250,-35)?

OpenStudy (anonymous):

Sorry (1250, 35)

OpenStudy (solomonzelman):

not k,h the other way around

OpenStudy (solomonzelman):

my fault

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y(x)=a(x-h)^2+k }\) will have a vertex of \(\color{#000000 }{ \displaystyle (h,k) }\)

OpenStudy (solomonzelman):

that is the form

OpenStudy (anonymous):

(35, 1250)

OpenStudy (solomonzelman):

yes, correct, and I apologize for my mistake.

OpenStudy (anonymous):

No worries, you have been amazing help!

OpenStudy (anonymous):

And guidance

OpenStudy (solomonzelman):

And the vertex of (35,1225) indicates that the maximum profit is 1225, and it occurs when you sell 35 dollars per item.

OpenStudy (solomonzelman):

Which statement best describes the intervals where the company's profit increases, decreases, or records a maximum? a) It is least when the selling price is $30. b) It is greatest when the selling price is $30. c) It decreases when the selling price increases from $10 to $35. d) It increases when the selling price increases from $10 to $35.

OpenStudy (anonymous):

d?

OpenStudy (solomonzelman):

you are close.

OpenStudy (anonymous):

c?

OpenStudy (solomonzelman):

can you sketch (somehow) a parabola that opens down please?

OpenStudy (solomonzelman):

any picture that looks like opening-down parabola would suffice.

OpenStudy (anonymous):